OpenStudy (anonymous):

Simplify. Write the answer in positive exponents. -(12)^0 +(-6)^0 Is the answer 2?

OpenStudy (anonymous):

its 0

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

-(12)^0 = -(1) and (-6)^0=1

OpenStudy (angela210793):

no it is 0 cause -(x)^0=-1 so -1+1=0

Directrix (directrix):

- 1 * (12)^0 + (-6)^0= -1* (12)^0 = -1 + 1 = 0

Directrix (directrix):

(12)^0 +(-6)^0 = 2